A 4.3kg mass is suspended by two strings, as shown on the right. If the tension in strong 1 is 28N, then determine the tension in string 2, and the angle.


1
Expert's answer
2021-03-14T19:20:29-0400
"W=T_2\sin{a}\\ T_2\cos{a}=T_1\\a=\arctan{\frac{W}{T_1}}=\arctan{\frac{4.3\cdot9.8}{28}}=56.4\degree"

"T_2\cos{56.4}=28\\T_2=50.6\ N"